OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of -1/2

OpenStudy (texaschic101):

(- 6, 2)....x = - 6 and y = 2 slope(m) = - 1/2 y = mx + b now we sub...we have everything but b....can u find b

OpenStudy (texaschic101):

2 = -1/2(- 6) + b now find b

OpenStudy (texaschic101):

ur answer will be y = -1/2x + (whatever u get for b)

OpenStudy (anonymous):

the answers I have to choose from is a. y= x/2-2 b. y=x/2-1 c. y=2x+1 d. y=-2x-1

OpenStudy (kayders1997):

First you could put it in point slope form that might be easier for you

OpenStudy (kayders1997):

Y-y1=m(x-x1) where y1 and x1 are the point they give you

OpenStudy (kayders1997):

So let's plug in the given information y-2=-1/2(x+6)

OpenStudy (anonymous):

okay now how do I solve that?

OpenStudy (kayders1997):

Distribute the -1/2

OpenStudy (kayders1997):

First so -1/2 times x is -1/2x now what is -1/2 times 6

OpenStudy (anonymous):

-3

OpenStudy (kayders1997):

Good so we have y-2=-1/2x-3

OpenStudy (kayders1997):

We need to have the y all by itself so what do you have to do?

OpenStudy (anonymous):

add 2

OpenStudy (kayders1997):

Yep :)

OpenStudy (kayders1997):

So you get?

OpenStudy (anonymous):

y=-1/2x-1

OpenStudy (kayders1997):

Yes :)

OpenStudy (anonymous):

so my answer would be d. -2x-1?